40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4325     Entry time: Fri Feb 18 17:52:25 2011
Author: josephb, valera 
Type: Update 
Category: CDS 
Subject: c1ass updated 

We updated the c1ass model to include the BS.  We removed the dither excitation of the PZTs.  PZT control goes to epics. To do this, modified the /cvs/cds/caltech/target/c1iscaux/PZT_AI.db file.  We basically have it sum both the existing epics slider and our new output from c1ass.

More importantly we updated the color scheme.

We compiled and tested the Dolphin and RFM which work.

I should note we can't figure out why testpoints are not working properly with just this model.  Alex and Joe spent well over an hour trying to debug it to no success.  Current workaround is to add what channels you want from c1ass to the DAQ recording.  Other testpoints on other models appear to be working.

Attachment 1: c1ass_updated.png  109 kB  | Hide | Hide all
c1ass_updated.png
ELOG V3.1.3-